Opened 10 years ago

Closed 3 years ago

#273 closed flight processing (fixed)

EU09/06, flight day 224b/2009, Vatna_01H

Reported by: mark1 Owned by: asm
Priority: alpha 4 medium Milestone:
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by mark1)

Data location: ~arsf/arsf_data/2009/flight_data/other/EU09_06-2009_224b_Vatna_01H

Data arrived from ARSF via SATA disk L on 7th October 2009.

Scientific objective:Glacier response to a changing climate in southern Iceland

Priority: alpha-4-medium

PI: David Graham

Sensors:

  • Eagle (13/05/2010)
  • Leica LIDAR (13/05/2010)

Change History (15)

comment:1 Changed 10 years ago by mark1

Still awaiting logsheet. No hawk data in directory so check logsheet when it arrives and check if there was a problem.

comment:2 Changed 10 years ago by mark1

  • Priority changed from immediate to alpha 4 medium

comment:3 Changed 10 years ago by chrfi

applanix navigation processing complete

note: supplied base station data did not correspond to flight time, downloaded new base station data through applanix

comment:4 Changed 10 years ago by emca

  • Owner set to emca
  • Status changed from new to assigned

Starting both LIDAR and Hyperspectral processing

comment:5 Changed 10 years ago by emca

Renamed the raw Eagle files to the VNIR* convention

comment:6 Changed 10 years ago by emca

Added PI and Proj Name to logsheet

comment:7 Changed 10 years ago by emca

LIDAR and Hyperspectral processing complete and ready for delivery check at:

~airborne/workspace/EU09_06-2009_224b_Vatna_01H/delivery/20100505/EU09_06
~airborne/workspace/EU09_06-2009_224b_Vatna_01H/delivery/20100506/EU09_06

comment:8 Changed 10 years ago by mark1

LIDAR delivery checked and looks good. Ready to deliver.

comment:9 Changed 10 years ago by mark1

Hyperspectral delivery checked. Example commands tested and work OK, lev1 data viewed and looks good. Ready to deliver.

comment:10 Changed 10 years ago by mark1

  • Description modified (diff)

Data dispatched to PI on 13/05/2010

comment:11 Changed 10 years ago by emca

  • Component changed from Processing: general to Archiving

All data has been delivered. Rsync'd to
/users/rsg/arsf/arsf_data/2009/flight_data/other/EU09_06-2009_224b_Vatna_01H

Ready for archiving

comment:12 Changed 3 years ago by lah

Archiving

Tidied and ready for upload to NEODC.

comment:13 Changed 3 years ago by asm

Archiving

Archived and it is available from CEDA. Processed data at http://browse.ceda.ac.uk/browse/neodc/arsf/2009/EU09_06/EU09_06-2009_224b_Vatna_01H/hyperspectral-reprocessed

Plymouth 12th May 2017

comment:14 Changed 3 years ago by asm

  • Owner changed from emca to asm
  • Status changed from assigned to new

comment:15 Changed 3 years ago by asm

  • Resolution set to fixed
  • Status changed from new to closed
Note: See TracTickets for help on using tickets.